OpenStudy (anonymous):

check out my answer! Multiply. (x + 6)(x − 6)

OpenStudy (anonymous):

FOIL First x times x = squared Outside -6timesx= -6x Inside 6 times x= 6x Last 6 times -6 = -36 = x squared - 6x+6x-36 +x squared -36 (-6x and 6x cancel out)

OpenStudy (anonymous):

x^2-36

OpenStudy (anonymous):

you got it. You can recognize that what you have is just the factored form of a perfect square.

OpenStudy (anonymous):

difference of 2 perfect squares

OpenStudy (anonymous):

(x+y)(x-y)=x^2-y^2 (x + 6)(x − 6)=x^2-36